# Lagrangian neural networks for nonholonomic mechanics

### v1

@@ 1 Introduction
In many cases, even when Euler-Lagrange equations are available, the solutions are not provided in analytical or explicit forms. Therefore, we can use numerical integrators to estimate the trajectories of a mechanical system. However, numerical integrators can sometimes produce poorly behaved trajectories concerning certain physical observables, such as energy. As an alternative, geometric integrators can be employed, since they are known to preserve energy (see, for instance, [2]). However, they may not be very accurate over long periods. Even worse is the case in which an analytical expression for the Lagrangian function is unknown or difficult to work with because we do not have a system of equations to solve.
@@
In recent years, there has been an increasing interest in using neural networks to address different issues of mechanical systems (see for example [7],[8],[11],[12],[14]). In this line, Lagrangian Neural Networks were introduced in [5] as an enhancement over other types of neural networks used in mechanical systems that do not preserve physical laws, providing a tool for scenarios where, for example, equations of motion are not available to get the actual trajectory. This method assumes that the Lagrangian of a mechanical system, a scalar function, can be parametrized using a neural network and be learned directly from the system’s data. That is, the goal of LNNs is to predict the Lagrangian function of a system based on data about its positions and velocities. This approach aims to represent the system’s equations of motion with a neural network while ensuring the preservation of some specific physical properties.
@@ 2 Lagrangian mechanics
The dynamics of a Lagrangian mechanical system is determined by Hamilton’s Principle, which states that a curve $q:[t_{0},t_{1}]\rightarrow Q$ is a trajectory of $(Q,L)$ if $q$ is a critical point of $S$ for infinitesimal variations $\delta q$ of $q$ with vanishing endpoints; that is, $dS[q]=0$ for all infinitesimal variations $\delta q$ of $q$ with vanishing endpoints.
This principle gives rise to a set of equations known as the Euler-Lagrange equations for the system $(Q,L)$. Thus, given a set of generalized coordinates $q=(q^{i})$ of the configuration space $Q$, a curve $q:[t_{0},t_{1}]\rightarrow Q$ is a solution of $(Q,L)$ if and only if
@@ 3 Lagrangian neural networks
As Cranmer et al. showed in [5], we can derive numerical expressions for the dynamics of Lagrangian systems by rewriting Euler-Lagrange equations in vectorized form and expanding the derivatives of the black-box Lagrangian to find the dynamics. Specifically, if we denote the generalized coordinates and velocities by $(q,\dot{q})$, we can write Euler-Lagrange equations (1) as follows:
$\frac{\mathrm{d}}{\mathrm{d}t}\nabla_{\dot{q}}L=\nabla_{q}L$ (2)
where $\ (\nabla_{\dot{q}})_{i}\equiv\frac{\partial}{\partial\dot{q}^{i}}$ and $\ \displaystyle(\nabla_{q})_{i}\equiv\frac{\partial}{\partial q^{i}}$. Expanding the time derivative, we obtain the expression:
@@ 4 Adding constraints
The application of LNNs to Lagrangian systems with holonomic constraints has been discussed in [7] and [15]. In this work, we will focus on nonholonomic constraints, which are restrictions that depend on both positions and velocities in a non-trivial manner, meaning they cannot be expressed solely in terms of positions, as is the case with holonomic constraints.
@@ 4.1 The vakonomic method
This is a set of ordinary differential equations of second order in positions and first order in Lagrange multipliers, which depends on $\Phi$ and $L$. Observe that LNNs could still be utilized for the augmented Lagrangian $\mathcal{L}$, as this can be seen as an unconstrained system for that Lagrangian. Nevertheless, from the point of view of implementation, it may be difficult to collect data about Lagrange multipliers to train a LNN model. This case requires special treatment and could be considered in a future work.
@@ 4.2 The nonholonomic method
In the nonholonomic approach, the system of differential equations describing the dynamics of the system may be also derived from a variational-like principle (known as Lagrange-d’Alembert principle) as follows. A curve $q(t)$ is a trajectory of the nonholonomic system if and only if the following equations are satisfied (summation over repeated indices is assumed from now on):
@@
By expanding the total time derivative, we obtain:
@@
Or equivalently,
$\frac{\partial^{2}L}{\partial\dot{q}^{j}\partial\dot{q}^{i}}\ddot{q}^{j}=\frac{\partial L}{\partial q^{i}}+\lambda_{a}\frac{\partial\Phi^{a}}{\partial\dot{q}^{i}}-\frac{\partial^{2}L}{\partial q^{j}\partial\dot{q}^{i}}\dot{q}^{j}$ (6)
@@ 4.3 Nonholonomic Lagrangian neural networks
For regular Lagrangians, the formulation above enables us to express the accelerations $\ddot{q}$ in terms of the Lagrangian $L$, the constraints $\Phi^{a}$, and the Lagrange multipliers $\lambda_{a}$, analogously to the unconstrained case, as follows.
@@
This equation can be seen as a generalization of equations obtained in [7] and [9] to the case where constraints depend on velocities in a non trivial manner.
@@
In the three examples we implement, the constraint happens to be linear, so we dedicate the next subsection to obtain a simpler expression of the previous equations for this type of restriction.
@@ 4.3.1 Linear constraints
As a result, we have the accelerations written as
@@ 5 Implementation
We evaluate our proposed method by applying it to three representative systems: a nonholonomic particle, a rolling wheel and a drone chasing a target. All these examples present linear nonholonomic constraints, so we follow the calculations described in Section 4.3.1. We implement the neural networks similarly to the approach developed in [5], as detailed in the subsequent subsection.
@@ 5.1 Datasets
Data for the systems of all examples are generated using 500 initial states $(q,\dot{q})$ satisfying the constraints, where positions are sampled using a uniform distribution of width 1 centered at 0 for cartesian coordinates, and a uniform distribution of width $\pi$ centered at 0 for angular coordinates. Then, from each one of these initial states, we numerically simulate 500 thousand-timestep constraint-compliant trajectories of the system for 15 time units (using an implementation of adaptive-step Dormand-Prince method) to give a total 500,000 data.
5.2 Training details
@@
For each model, we use a four-layer architecture with two hidden layers, each containing 128 fully connected neurons, and an output layer with a single neuron. The first layer’s configuration depends on the specific example. We employ softplus activations and use adaptive learning rates starting at $10^{-3}$. In each example, we follow a stochastic gradient descent strategy, training the models for 300 epochs. We take minibatches of size 1000, so that each minibatch corresponds to a complete trajectory of the dataset.
We have used 88% of the total data generated for training and we reserved 12% for testing (that is the last 60 trajectories).
@@ 5.3 Performance
In each example, we evaluate the performance of both models by comparing the dynamics obtained from the prediction of the Lagrangian learned from LNN and from LNN-nh models. Loss function is taken to be the mean squared error in accelerations. For energy and constraint assessment, we simulate 5 different trajectories in the same way we did for training and testing set and compute energy and constraint along these trajectories.
@@ 6.1 The nonholonomic particle
Consider a free particle of mass 1 moving in three-dimensional space with standard coordinates in $\mathbb{R}^{3}$, subjected to the nonholonomic constraint given by $\Phi(q,\dot{q})=\Phi(x,y,z,\dot{x},\dot{y},\dot{z})=\dot{z}-y\dot{x}=0$. Notice that the constraint $\Phi$ is linear in velocities since $\Phi(q,\dot{q})=\omega(q)\cdot\dot{q}$ with the differential 1-form $\omega(q)=-ydx+dz,$ that is $(-y,0,1)$ in coordinates.
The Lagrangian of the system is given by $\ L(q,\dot{q})=L(x,y,z,\dot{x},\dot{y},\dot{z})=\displaystyle\frac{1}{2}(\dot{x}^{2}+\dot{y}^{2}+\dot{z}^{2})$ and then, following Section 4.3.1, we have $\ M=1+y^{2}$ and $\ f=(0,0,0).$ Subsequently, we obtain the equation of motion of the nonholonomic particle given by
@@
Figure 1 compares the learned and true value of each coordinate acceleration in the example for both models, exhibiting a major dispersion in LNN learned accelerations.
Figure 1: True vs. learned cartesian accelerations for LNN-nh and LNN models in the nonholonomic particle example.
@@
6.2 Drone chasing a target
Consider a target and a drone moving in the plane. We take the target as a particle of mass $m_{t}$ moving freely along the $y$ axis and we take the drone as another particle of mass $m_{d}$ moving in the plane chasing the target, i.e. with velocity pointing directly to it (see [13] for details). The position of the target is determined with a single coordinate $w$, whereas position of the drone may be described with two cartesian coordinates $(x,y)$. Hence, the state of the system is completeley described by the tuple $(q,\dot{q})=(w,x,y,\dot{w},\dot{x},\dot{y})$.
@@ 6.2 Drone chasing a target
Figure 3: True vs. learned cartesian accelerations for LNN-nh and LNN models in the target-drone example.
@@
Figure 4: (4(a)) Comparison of the loss function from the target-drone example for training and testing sets corresponding to LNN and LNN-nh models. Picture (4(b)) shows the evolution of the constraint function along five trajectories of the drone-target system generated from the same initial conditions for both models. Picture (4(c)) exhibits the total energy of each trajectory normalized with the corresponding constant true energy. Picture (4(d)) shows the mean relative error in energy of the trajectories.
@@ 6.3 A vertical rolling wheel
Figure 5: True vs. learned generalized accelerations for LNN-nh and LNN models in the wheel example.

### v2

@@ 1 Introduction
In many cases, even when the Euler–Lagrange equations are available, the solutions are not given in analytical or explicit form. In such cases, numerical integrators can be used to estimate the trajectories of a mechanical system. However, these methods may yield trajectories that behave poorly with respect to certain physical observables, such as energy. As an alternative, geometric integrators can be employed, as they are known to preserve energy (see, for instance, [8]). Nevertheless, their accuracy may degrade over long time intervals. An even more challenging situation arises when an analytical expression for the Lagrangian is unavailable or difficult to handle, leaving us without a system of equations to solve.
In the presence of constraints, mechanical systems can be broadly classified as holonomic or nonholonomic, depending on the nature of the constraints. Holonomic constraints can be expressed as algebraic equations involving only the generalized coordinates. These constraints are integrable and reduce the dimension of the configuration space; the dynamics of such systems can be derived from a classical variational principle by incorporating the constraints through Lagrange multipliers. This framework preserves key geometric properties, such as symplecticity and, under suitable conditions, energy conservation.
In contrast, nonholonomic systems are characterized by non-integrable, velocity-dependent constraints, and differ fundamentally from holonomic systems both in their geometric structure and in the formulation of their equations of motion. As a result, the study of nonholonomic mechanical systems is a classical subject that has remained highly active over the past few decades, driven by both practical applications—such as robotics, locomotion, and control—and by questions of intrinsic mathematical interest. Representative works in the field include [7, 10, 16, 18, 28, 25, 5, 3, 4, 6, 27, 14, 12].
Unlike holonomic constraints, which restrict the configuration space, nonholonomic constraints define particular admissible velocities. As a consequence, the dynamics of nonholonomic systems are not derived from Hamilton’s principle but rather from the Lagrange-d’Alembert principle (when the constraint is linear in velocities) or the Chetaev principle (for generic constraints, not necessarily linear in velocities), and they do not, in general, preserve symplectic structure or energy. This intrinsic difference poses serious challenges for standard numerical integrators, which often fail to preserve the constraints or yield long-term energy drift. In fact, many classical integration schemes do not respect the geometry of nonholonomic flows, leading to qualitative inaccuracies in the simulation of such systems.
@@
In recent years, there has been an increasing interest in using neural networks to address different issues of mechanical systems (see for example [20],[21],[24],[26],[30]). In this line, Lagrangian Neural Networks were introduced in [13] as an enhancement over other types of neural networks used in mechanical systems that do not preserve physical laws, providing a tool for scenarios where, for example, equations of motion are not available to get the actual trajectory. This method assumes that the Lagrangian of a mechanical system, a scalar function, can be parametrized using a neural network and be learned directly from the system’s data. That is, the goal of LNNs is to predict the Lagrangian function of a system based on data about its positions and velocities. These models incorporate physical structure into the learning process and have shown improved performance over standard integrators in preserving key properties such as energy. This approach aims to solve the system’s dynamics from a Lagrangian learned by a neural network while ensuring the preservation of some specific physical properties. Given these advantages, it is natural to explore whether a properly adapted version of LNNs can offer similar benefits in the nonholonomic setting, where traditional methods struggle.
From a theoretical standpoint, it is well known that the same laws do not always govern the dynamics of constrained systems as those of unconstrained systems. Since LNNs, as introduced in [13], are trained using the equations of motion for unconstrained systems, it is natural to ask whether such models can still learn meaningful Lagrangian functions when applied to systems with constraints. Some previous works, such as [20] and [32], have addressed holonomic constraints.
The motivation for this work is to investigate whether a targeted adaptation of LNNs can more effectively handle nonholonomic constraints. As shown in the examples we present, standard LNNs fail to preserve the constraint along the trajectories generated by the model, even when starting from initial states that satisfy the constraint.
To address this issue, we propose modifying the loss function of the Lagrangian neural network to explicitly account for nonholonomic constraints during training, when such constraints are present in the mechanical system. In the examples studied, this modification yields in solutions that more closely adhere to the system’s physical properties -particularly in terms of energy conservation for most of the examples considered, and most notably in the preservation of the nonholonomic constraint. By doing so, we aim to extend the applicability of Lagrangian-based learning to this important and challenging class of mechanical systems.
@@ 2 Lagrangian mechanics
[…]
@@ 3 Lagrangian neural networks
As explained in the previous section, the standard modeling of mechanical systems relies on the assumption that the Lagrangian function $L$ is explicitly known. From this expression, one can symbolically derive the Euler–Lagrange equations to obtain the system’s equations of motion. However, as demonstrated by Cranmer and collaborators in [13], it is possible to treat $L$ as a black box, thereby allowing the modeling of the system’s dynamics without requiring an analytical expression for the Lagrangian or its associated equations of motion. Within this framework, the Lagrangian formalism still permits a numerical approximation of the system’s dynamics. By rewriting the Euler–Lagrange equations in vectorized form and solving for the accelerations, one can compute $\ddot{q}$ from input data $(q,\dot{q})$, even when the Lagrangian is implicitly represented by a neural network. This approach forms the basis for data-driven methods in Lagrangian mechanics, enabling the estimation of physically consistent dynamics directly from observed trajectories.
In [13], the authors proposed to derive numerical expressions for the dynamics of Lagrangian systems and expanding the derivatives of the black-box Lagrangian to find the dynamics. Specifically, if we denote the generalized coordinates and velocities by $(q,\dot{q})$, we can write Euler-Lagrange equations (1) in vectorized form as follows:
$\frac{\mathrm{d}}{\mathrm{d}t}\nabla_{\dot{q}}L=\nabla_{q}L.$ (2)
where $\ (\nabla_{\dot{q}})_{i}=\frac{\partial}{\partial\dot{q}^{i}}$ and $\ \displaystyle(\nabla_{q})_{i}=\frac{\partial}{\partial q^{i}}$ are the components of vectors $\nabla_{\dot{q}}$ and $\nabla_{q}$, respectively. Expanding the time derivative and denoting $\nabla_{\dot{q}}^{T}$ the transpose of the vector $\nabla_{\dot{q}}$, we obtain the expression:
@@
Thus, given a data set of coordinate pairs $(q,\dot{q})$, we have a procedure for computing the accelerations $\ddot{q}$ from a black‑box Lagrangian. Using the chain rule for backpropagation process, we take the required derivatives of the network‑estimated Lagrangian and solve the vectorized Euler–Lagrange equations for the accelerations according to equation (4). We then define a loss function as the mean‑squared error (MSE) between these predicted accelerations and the ground‑truth accelerations in the training data. By minimizing this loss, the network learns a Lagrangian that best reproduces the observed dynamics.
@@ 4 Adding constraints
Lagrangian neural networks have been successfully applied to a wide range of mechanical systems, including unconstrained systems, systems with holonomic constraints (as discussed in [20] and [32]), and even externally actuated systems, such as in [19] and [31]. In contrast, our work focuses on systems without external forcing, but with nonholonomic constraints. In this paper, we develop a different extension of the LNN framework by incorporating nonholonomic constraints directly into the learning process. These constraints depend on both positions and velocities in a nontrivial way, meaning they cannot be expressed solely in terms of positions, as is the case with holonomic constraints.
@@ 4.1 The vakonomic method
This is a set of ordinary differential equations of second order in positions and first order in Lagrange multipliers, which depends on $\Phi$ and $L$. Observe that LNNs could still be utilized for the augmented Lagrangian $\mathcal{L}$, as this can be seen as an unconstrained system for that Lagrangian. However, augmented Lagrangians are always singular, a kind of Lagrangian that was not treated in the examples considered in [13]. On the other hand and from the point of view of implementation, it may be difficult to collect data about Lagrange multipliers to train a LNN model. For these reasons, the case of vakonomic constraints requires special treatment and could be considered in a future work.
@@ 4.2 The nonholonomic method
In the nonholonomic approach, the system of differential equations describing the dynamics of the system may be also derived from a variational-like principle as follows. A curve $q(t)$ is a trajectory of the nonholonomic system if and only if the following equations are satisfied (summation over repeated indices is assumed from now on):
@@
By expanding the total time derivative, we obtain the equations of motion as follows:
@@
or, equivalently,
$\frac{\partial^{2}L}{\partial\dot{q}^{j}\partial\dot{q}^{i}}\ddot{q}^{j}=\frac{\partial L}{\partial q^{i}}+\lambda_{a}\frac{\partial\Phi^{a}}{\partial\dot{q}^{i}}-\frac{\partial^{2}L}{\partial q^{j}\partial\dot{q}^{i}}\dot{q}^{j}\ .$ (6)
@@ 4.3 Nonholonomic Lagrangian neural networks
We recall that the central question of this work is whether a tailored adaptation of Lagrangian neural networks can more effectively accommodate nonholonomic constraints. To address this, we propose modifying the procedure used to compute accelerations during the learning process of the LNN so as to explicitly incorporate the constraints when present in the mechanical system. The loss function is then defined using the appropriate expression for the accelerations dictated by nonholonomic dynamics.
For the purpose of learning a regular Lagrangian, and assuming the constraints are known, the above formulation allows us to express the accelerations $\ddot{q}$ in terms of the Lagrangian $L$, the constraints $\Phi^{a}$, and the Lagrange multipliers $\lambda_{a}$, in a manner analogous to the unconstrained case, as follows.
@@
This equation can be seen as a matrix version of equations in [17] and a generalization of equations obtained in [20] and [22] to the case where constraints depend on velocities in a non trivial manner.
@@
In three of the examples we implement, the constraint happens to be linear, so we dedicate the next subsection to obtain a simpler expression of the previous equations for this type of restriction.
@@ 4.3.1 Linear constraints
As a result, we have the accelerations for a system with linear nonholonomic constraints written as
@@ 5 Implementation
We evaluate our proposed method by applying it to four representative systems: a nonholonomic particle, a rolling wheel, a dog chasing a man and a particle subjected to a constant gravitational field. The first three examples present linear nonholonomic constraints, so we follow the calculations described in Section 4.3.1 for these cases. We implement the neural networks similarly to the approach developed in [13], as detailed in the subsequent subsection.
@@ 5.1 Datasets
We generate data for each system by taking 500 true trajectories as follows: first, we sampled 500 initial states of the form $(q,\dot{q})$ fulfilling the constraint. The initial states for all examples were sampled using an uniform distribution on a fixed range (minimum and maximum values of the range depend on the specific example). Then, we simulate 500 different trajectories starting at the mentioned initial states for 1000 timesteps (ranging from 0 to 10) using the true equations of motion (that is, the equations derived analytically using the true known Lagrangian of the system). Thus, we end up with 500,000 pairs of the form $(q,\dot{q})$.
This simulation was made with an implementation of an adaptive-step Dormand-Prince method in the examples with linear constraints, so it is a numerical simulation, but we refer to these trajectories as true trajectories because they were generated using the true Lagrangian. In the example of the particle subjected to a constant gravitational field, equations can be solved analytically, so we were able to simulate trajectories without the use of any numerical integrator.
Finally, for every pair $(q,\dot{q})$ in each dataset, acceleration targets $\ddot{q}_{true}$ are generated using the true Lagrangian and the constraint function together with Eq. (8), for LNN-nh model, and the true Lagrangian together with Eq. (4) for LNN model.
5.2 Training details and architecture
@@
For each model (LNN and LNN-nh), we use a four-layer MLP architecture. The neural network will be trained to learn the Lagrangian function of each example. Thus the output layer will have a single unit and the input layer must have as many units as the number of positions and velocities of the system. We choose the two hidden layers to contain 128 fully connected neurons each. We employ softplus for activations and use ADAM optimizer with adaptive learning rates starting at $10^{-3}$. In each example, we follow a stochastic gradient descent strategy, training the models for 300 epochs. During each epoch we generate a dataset of 500 true trajectories fulfilling the constraint (as detailed in the previous subsection). We then take minibatches of size 1000, so that each minibatch corresponds to a complete trajectory of the dataset generated in the current epoch. We divide the dataset of each epoch, using 88% for training and the remaining 12% for testing (that is the last 60 trajectories of every epoch).
All the implementation was carried out using JAX libraries  jax.example_libraries.stax  and
jax.example_libraries.optimizers.
@@ 5.3 Performance
In each example, we evaluate the performance of both models by considering the dynamics obtained from the prediction of the Lagrangian learned from LNN and from LNN-nh models. We compare loss function, predictions, energy and constraint from both models.
Loss function is taken to be the mean squared error in accelerations, as said before. More specifically, for each pair $(q,\dot{q})$ and corresponding target $\ddot{q}_{true}$ we compute the predicted $\ddot{q}_{pred}$ using the Lagrangian learned with the LNN-nh model and the constraint together with Eq. (8), and the Lagrangian learned with the LNN model together with Eq. (4). Then we take the MSE of the difference $\ddot{q}_{true}-\ddot{q}_{pred}$ and sum and normalize over all the minibatch.
For energy and constraint assessment, we simulate five different trajectories for each model (ten trajectories total) following the process outlined in subsection 5.1 but this time using the Lagrangians learned by the models instead of the true Lagrangian (we call these trajectories learned trajectories to distinguish them from the true trajectories). We stress that these trajectories are not part of the training nor the testing set, since they are simulated using the models. We then compute the true energy and constraint along these learned trajectories.
@@ 6 Examples
In this section, we illustrate the performance of our proposed method across four representative mechanical systems subjected to nonholonomic constraints. The first three examples involve systems with linear velocity constraints, while the fourth example presents a nonlinear nonholonomic constraint. These examples illustrate how the proposed adaptation of Lagrangian neural networks performs when learning the dynamics of systems subjected to different types of nonholonomic constraints. For each system, we compare the LNN-nh model with a standard LNN by analyzing the loss function, the predicted accelerations and evaluating how well each approach preserves energy and satisfies the constraints along learned trajectories. The results suggest that incorporating the constraint structure during training can lead to improved performance in capturing the system’s dynamics.
@@ 6.1 The nonholonomic particle
[…]
@@
For the example, we randomly selected five hundred pairs $(q,\dot{q})$ and computed the corresponding accelerations using both models. The resulting scatter plots, shown in Figure 1, compare each component of $\ \ddot{q}_{pred}$ and $\ \ddot{q}_{true}$ for the LNN and LNN-nh models. The plots reveal a greater dispersion in the accelerations predicted by the LNN model.
@@
Figure 1: The scatter plots show five hundred true vs. corresponding learned cartesian accelerations for LNN-nh and LNN models in the nonholonomic particle example.
@@
6.2 A dog pursuing a man
[…]
@@ 6.2 A dog pursuing a man
Figure 3: The scatter plots show five hundred true vs. corresponding learned cartesian accelerations for LNN-nh and LNN models in the man-dog example.
@@
Figure 4: (4(a)) Comparison of the loss function from the man-dog example for training and testing sets corresponding to LNN and LNN-nh models. Picture (4(b)) shows the evolution of the constraint function along five trajectories of the man-dog system generated from the same initial conditions for both models. Picture (4(c)) exhibits the total energy of each trajectory normalized with the corresponding constant true energy. Picture (4(d)) shows the mean relative error in energy of the trajectories.
@@ 6.3 A vertical rolling wheel
Figure 5: The scatter plots show five hundred true vs. corresponding learned accelerations for LNN-nh and LNN models in the wheel example.
@@
6.4 A system with a nonlinear nonholonomic constraint: Appell’s example
Consider a point particle of mass $m$ moving in space subjected to a constant gravitational field of strength $g$. A configuration is then given by $(x,y,z)\in\mathbb{R}^{3}$. The Lagrangian of this system may be written as
$L(q,\dot{q})=L(x,y,z,\dot{x},\dot{y},\dot{z})=\displaystyle\frac{m}{2}(\dot{x}^{2}+\dot{y}^{2}+\dot{z}^{2})-mgz,$
In addition, the particle is constrained to move according to the nonholonomic constraint
$\Phi(x,y,z,\dot{x},\dot{y},\dot{z})=b^{2}(\dot{x}^{2}+\dot{y}^{2})-\dot{z}^{2}=0,$
for some constant $b>0$. See for instance [17, 1] and references therein for a more detailed analysis of the system.
Unlike the linear examples, the force $f$ is now nonzero and given by $f=(0,0,-mg)$. In this case the matrix $M$ is scalar and given by
$M=\frac{1}{m}(4b^{4}\dot{x}^{2}+4b^{4}\dot{y}^{2}+4\dot{z}^{2}).$
The final equations of motion obtained by using Chetaev’s principle and eliminating the Lagrange multiplier as we did in previous examples are
$\begin{pmatrix}\ddot{x}\\ \ddot{y}\\ \ddot{z}\end{pmatrix}=\frac{-b^{2}g}{b^{4}\dot{x}^{2}+b^{4}\dot{y}^{2}+\dot{z}^{2}}\begin{pmatrix}\dot{x}\dot{z}\\ \dot{y}\dot{z}\\ b^{2}\dot{x}^{2}+b^{2}\dot{y}^{2}\end{pmatrix}.$
Figure 7 compares the learned and true value of each coordinate acceleration for both models. While both models exhibit reasonable predictions, we can see that accelerations predicted by the LNN model show a slight bias. Figure 8, in turn, shows the evolution of loss functions during training for testing and training sets of the LNN and LNN-nh models. The same figure also displays the performance of the learned trajectories using both systems of the energy and constraint functions. All computations were performed with parameters $m=1$ and $b=1$.
Figure 7: The scatter plots show five hundred true vs. corresponding learned accelerations for LNN-nh and LNN models in Appell’s example.
(a) Loss function
(b) Nonholonomic constraint
(c) Total energy
(d) Mean error of energies
Figure 8: (8(a)) Comparison of the loss function from Appell’s example for training and testing sets corresponding to LNN and LNN-nh models. Picture (8(b)) shows the evolution of the constraint function along five trajectories of the nonholonomic particle generated from the same initial conditions, first using the Lagrangian learned from a LNN model and second using a LNN-nh model. Picture (8(c)) exhibits the total energy of each trajectory normalized with the corresponding constant true energy. In Picture (8(d)) can be seen the mean relative error in energy of the trajectories.